Circular No. 7959

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2002ep, 2002eq, 2002er, 2002es, 2002et The discoveries of several apparent new supernovae have been reported

from unfiltered CCD frames. SNe 2002ep and 2002eq were reported by W. M. Wood-Vasey et al. from Palomar NEAT images (cf. IAUC 7953), while the remaining three objects were found on LOTOSS KAIT images (W. D. Li, B. Swift, and M. Ganeshalingam; cf. IAUC 7906).

SN  2002 UT  α_{2000}  δ_{2000}  Mag.  Offset
2002ep Aug. 6.47 21^{h}57^{m}21^{s}.66 − 7^{o}51′24′′.8 18.7 ∼ 5′′ W
2002eq Aug. 8.36 21 19 00.19 −16 17 12.7 19.6 ∼ 4′′ S
2002er Aug. 23.2 17 11 29.88 + 7 59 44.8 17.5: 12′′.3 W, 4′′.7 N
2002es Aug. 23.5 3 23 47.23 +40 33 53.5 16.4: 18′′.8 W, 25′′.5 N
2002et Aug. 24.3 20 08 31.02 −25 27 37.5 16.5: 2′′.1 W, 6′′.8 S

Additional magnitude estimates with the same instrumentation for each respective object: SN 2002ep, July 13 UT, [20.5; 23, [20.5; Aug. 7.5, 18.7; 17.4, 19.0. SN 2002eq, July 4, [20.5; 19, [20.5; 18.3, 19.4. SN 2002er in UGC 10743, Aug. 16.2, [19.0:, 24.2, 16.9:. SN 2002es in UGC 2708, Feb. 15.2, [19.0:; Aug. 12.5, 18.5:, 24.5, 16.3:. SN 2002et in MCG –04-47-10, Aug. 2.3, [19.0:, 10.3, hint; 25.3, 16.5:.

1997 CQ_{29}

K. Noll and D. Stephens, Space Telescope Science Institute; W. Grundy, J. Spencer, R. Millis, and M. Buie, Lowell Observatory; D. Cruikshank, Ames Research Center, NASA; S. Tegler, Northern Arizona University; and W. Romanishin, University of Oklahoma, report on confirming observations of the binary transneptunian object 1997 CQ_{29} (cf. IAUC 7824), obtained with the Hubble Space Telescope during June 18.312– 18.351, 30.131–30.170, and July 12.152–12.191 UT: “On each date, three 800-s exposures were obtained with the F814W filter (roughly I band), the target being placed on the PC chip (0′′.045 pixel scale) of the WFPC2 camera. The binary was easily detected on all three dates with a signal-to-noise ratio of 25 or more for each component. The components were separated by an average of 0′′.334 (≈ 8 pixels). The position angle measured from component A to component B was seen to change slowly from 334^{o} ± 1^{o} on June 18 to 340^{o}.7 ± 0^{o}.8 on July 12, confirming orbital motion.”

2002 August 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green